total fixed cost=Rm100 per period. average function is AVC=60-12Q+Q^2. AVC represents average variables cost(RM) Q represents rate of output(units)
A. determine TVC function
B. What is TC function?
C. Determine MC function?
D. Calculate the rate of output at which marginal cost is minimum?
E. compute min marginal cost?
1
Expert's answer
2020-05-19T09:55:50-0400
"AVC="

"TVC\over Q"


"TVC=Q\u00d7AVC"


"Q(60-12Q+Q^2"


"60Q-12Q^2+Q^3"



"TC =TFC+TVC"


"=Rm100+60Q-12Q^2+Q^3"


"MC ="

"d_{Tc}\over d_Q"


"60- 24Q+3Q^3"

D.

"d_{MC}\over d_ Q"

"=0"


"-24+6Q=0"


"6Q=24"


"Q=4"

E.

"60-24Q+30Q^2"


"60-24(4)+3(4)^2=12"